Sunspot Beverages, Ltd., of Fiji uses the weighted-average method in its process costing system. It makes blended tropical fruit drinks in two stages. Fruit juices are extracted from fresh fruits and then blended in the Blending Department. The blended juices are then bottled and packed for shipping in the Bottling Department. The following information pertains to the operations of the Blending Department for June.

Percent Completed
Units Materials Conversion
Work in process, beginning 100,000 70% 40%
Started into production 477,500
Completed and transferred out 467,500
Work in process, ending 110,000 75% 25%


Materials Conversion
Work in process, beginning $ 41,300 $ 15,100
Cost added during June $ 442,700 $ 321,500


Required:

1. Calculate the Blending Department's equivalent units of production for materials and conversion in June.

2. Calculate the Blending Department's cost per equivalent unit for materials and conversion in June.

3. Calculate the Blending Department's cost of ending work in process inventory for materials, conversion, and in total for June.

4. Calculate the Blending Department's cost of units transferred out to the Bottling Department for materials, conversion, and in total for June.

5. Prepare a cost reconciliation report for the Blending Department for June.
Answer #1

1)

MATERIAL CONVERSION
Units completed and transferred out 467500 467500
Ending work in process inventory 110000*75%=82500 110000*25%= 27500
Equivalent units 550000 495000

2)

MATERIAL CONVERSION
Work in process ,Beginning 41300 15100
Current period cost 442700 321500
Total cost to account for 484000 336600
Equivalent units 550000 495000
cost per equivalent unit 484000/550000=$ .88 336600/495000=.68

3)

Ending work in process
cost per equivalent unit Equivalent unit Total
Material .88 82500 72600
Conversion .68 27500 18700
Total 91300

4)

cost of units completed and transferred out
cost per equivalent unit Equivalent unit Total
Material .88 467500 411400
Conversion .68 467500 317900
Total 729300

5)

BLENDING DEPARTMENT

COST RECONCILIATION

Cost to account for
Beginning work in process(41300+15100) 56400
Current period cost (442700+321500) 764200
Total cost to account for 820600
cost accounted for
cost of units completed and transferred out 729300
cost of ending work in process 91300
Total cost accounted for 820600